Urine dipstick leukocyte esterase testing in a group of 1000women, 300 (30%) of whom have a UTI (as determined by a gold-standard test such as urine culture). The dipstick test has sensitivity of 71% and specificity of 85%.

How many women with a UTI would have a negative result?
Answer #1

Tp( True positive) = 300

Sn = tp/tp+fn 71/100= 300/300+fn fn= 122

Fn + tp= 422 , tn+ fp= 1000-422= 678

Sp= tn/tn+fp. 85/100= tn/678 = 576